# Preference for Verifiability

### v1

@@ 5 Axioms
This section provides a set of descriptive axioms that characterize expected verification and obfuscation utilities. Such axioms allow us to test whether decision makers’ behavior is consistent with our decision models.
We assume the existence of a biseparable utility representation. Biseparable preferences allow us to identify the decision weights of individuals with respect to binary acts while making minimal assumptions about their behavior. The axioms for biseparable utility are provided in [29] and for convenience restated in appendix G.
@@
Thus, in the biseparable model, decision makers’ preferences are only meaningfully restricted on binary acts. Under the assumption that certainty equivalents exist and preferences are biseparable, we can define preference averages of consequences and acts. These preference averages, first introduced in [30], allow us to use [22] style axioms without resorting to objective probabilities being available. While objective probability lotteries can easily be implemented in a laboratory setting, for applications in the context of greenwashing behavior, such lotteries are often unavailable.
@@
Suppose Trees and RECs refer to investing a fixed amount of money into the relevant carbon offset certificates. Suppose further that there are no returns to scale to either technology. If the decision maker is risk neutral (in terms of the curvature of $u$) over the amount of CO2 emissions reduced, then $1/2$ Trees $\oplus 1/2$ RECs refers to the act in which the carbon emissions in every state of the world are the arithmetic average of Trees and RECs and can be thought of as investing half the money into each technology. However, if the decision maker does not have a linear utility over the achieved reduction, this will generally not be the case. Instead, in each state $1/2$ Trees $\oplus 1/2$ RECs would offer the certainty equivalent of a (perfectly verifiable) equal probability lottery of the reduction achieved by Trees and RECs on that state. ∎
@@
Thus, two acts are comonotonic, if they agree on the ranking of states according to whether they achieve more desirable consequences. Two acts $a,b$ are not comonotonic if we can find two states $s,s^{\prime}$ such that according to $a$, state $s$ yields a strictly better consequence than $s^{\prime}$ and act $b$ yields a strictly worse consequence on $s$ than $s^{\prime}$. Notice that constant acts are comonotonic with all acts.
@@
If the acts $a,b,c$ are comonotonic then for every verifiable event the state that achieves the worst consequence is preserved by pointwise preference averages.
In the context of preference for verifiability, comonotonic independence has a natural interpretation. If we assume that the decision maker cares about which consequences can be proven to have obtained, comonotonic independence emerges as a natural condition. Pointwise mixing two acts $a,b$ with a third $c$ may affect which consequences can be proven to have obtained. However, as long as all three acts are comonotonic, the effect will be symmetric and no preference reversal should occur.
@@
In our example, all acts are comonotonic and thus decision makers with an expected verification utility or obfuscation utility fulfill independence with respect to these three acts. However, once we also allow for acts that yield a uniquely best consequence on $u$ (or $s$) a decision maker with an expected verification or obfuscation utility would violate independence. Such acts could for example be the possibility to purchase an insurance on high prices of RECs which allow the decision maker to achieve a higher carbon reduction in case state $u$ obtains. ∎
Axiom 2 (Supermodularity).
@@
$1/2[{\gamma E\cup F\beta}]\oplus 1/2[{\gamma E\cap F\beta}]\quad\succsim(\precsim)\quad 1/2[{\gamma E\beta}]\oplus 1/2[{\gamma F\beta}].$ (4)
Supermodularity effectively states that given events $E$ and $F$, the value of getting a good consequence on $E\cup F$ consists at least of the value of getting a good consequence on $E$ and on $F$ net the value of getting a good consequence on $E\cap F$. It may be strictly greater in case there are verifiable events that are neither a subset of $E$ nor $F$ but are contained in $E\cup F$. In this case, the decision maker is able to prove on some additional states that a good consequence has been achieved. Submodularity states the reverse, the value of getting a good consequence on $E\cap F$ is at most the value of getting a good consequence on $E$ and $F$ net the value of the intersection.
@@
A firm trying to obfuscate that their CO2 compensation program is ineffective has an interest to choose a compensation program that on any ex-post verifiable event is at least in some state consistent with a high reduction of CO2. This generates submodularity since achieving a good consequence on $E$ may also be helpful if event $F$ is verifiable. If the stakeholder proves that event $F$ (or, $E$) arises, the firm can then point to a state in $E\cap F$ and argue that the subjective probability of this state is highly likely. ∎
Definition 5 (Min-increasing event).
Let $\gamma\succ\beta$. An event $E$ is min-increasing if $\gamma_{E}\beta\succ\gamma_{E-F}\beta$ for all nonnull events $F\subset E$.
In our verification utility, min-increasing events are the events $E$ for which the decision maker has a cover of verifiable events $V$ that are all subsets of $E$. On a min-increasing event, obtaining at least $\gamma$ on every subevent is critical or the utility will decrease. If an event $E$ is not critical, then we can make $\gamma_{E}\beta$ worse on some state and the decision maker is indifferent to such a change.
Definition 6 (Max-increasing event).
An event $E$ is max-increasing if $\beta_{E-F}\gamma\succ\beta_{E}\gamma$ for all nonnull events $F\subset E$ and some consequences $\gamma\succ\beta$.
An event $E$ is max-increasing if $\beta_{E\cup F}\gamma\succ\beta_{E}\gamma$ for all nonnull events $F\subset\overline{E}$ and some consequences $\gamma\succ\beta$.
In our obfuscation utility, max-increasing events are the events $E$ for which adding a better consequence on any state strictly increases the utility.
Max-increasing events play in obfuscation utility exactly the same role as min-increasing events play in verification utility. We therefore refer by the term critical events to either the min-increasing or the max-increasing events, depending on the context.
Axiom 3 (Critical event modularity).
[…]
  1. 1.
    $E\cap F$ is a critical event,
  2. 2.
    $E\cup F$ is a critical event, and
  3. 3.
    $1/2[{\gamma A\beta}]\oplus 1/2[{\gamma A\cap E\cap F\beta}]\sim 1/2[{\gamma A\cup E\beta}]\oplus 1/2[{\gamma A\cap F\beta}]$
The first two conditions are intuitive. For example, if $E$ and $F$ are verifiable, then their intersection is also verifiable (by providing both the proof that $E$ and the proof that $F$). Their union is then critical because no longer achieving a good consequence on every state of both events means that it is no longer possible on $E$ or $F$, or both, to prove that a good consequence has been obtained.
Example.
If $E$ and $F$ are critical, then the decision maker expects to be able to prove this in case these events actually obtain. Notice that in this case $E\cap F$ is also critical and so is $E\cup F$. Suppose for the moment that $A=E\cup F$. Critical event modularity states that the value obtained from obtaining a good consequence on both $E$ and $F$ can be separated into the value from obtaining a good consequence on $E$ and obtaining a good consequence on $F$. If $A$ is a subset of their union, then critical event modularity generalizes this separability even for non-critical $A\subset E\cup F$. ∎
5.1 Characterization of Expected Verification Utility
The above stated axioms contain necessary and sufficient conditions on preferences for an expected verification utility. More precisely, we obtain the following characterization result:
@@ 5.1 Characterization of Expected Verification Utility
Suppose $\succsim^{1}$ and $\succsim^{2}$ are expected verification utilities. Then $a\succsim^{1}b\Leftrightarrow a\succsim^{2}b$ for all $a,b\in\mathscr{A}$ if and only if:
@@
In our example, the verifiability of $\{s,t\}$ and $\{s,t,u\}$ can be inferred from behavior of the decision maker and need not be exogeneously given. Moreover, the subjective probabilities of $\{s,t\}$ and $\{u\}$ are uniquely determined from behavior. ∎
Effectively, if preferences over acts are identical then if one decision maker can prove that a certain event obtains, the other decision maker is able to prove that this event obtains or even able to prove that a strict subset of this event obtains. Thus, it could be the case that decision maker 1 is able to prove that $E,F,E\cup F,E\cap F$ obtain (whenever they actually obtain) while decision maker 2 is only able to prove that $E,F,E\cap F$ obtain. In terms of behavior, we would not see a difference because the decision maker will always want to provide the most precise proof possible.
@@ 5.2 Characterization of Expected Obfuscation Utility
We also obtain a dual theorem that characterizes obfuscation utility. Since an expected obfuscation utility is equivalent to an expected verification utility with a reverse order of preferences, supermodularity is replaced by submodularity and critical event modularity holds with respect to max-increasing instead of min-increasing events.
@@
Naturally, the expected obfuscation utility has the same uniqueness properties as expected verification utility. Expected obfuscation utility can be interpreted as the decision maker minimizing the utility of a stakeholder with expected verification utility who has a reverse order of preference of consequences. Similarly, expected verification utility can be seen as minimizing the utility of a stakeholder who wants to “find a hair in the soup”, i.e., who wants to point at the possibility that a bad consequence may have been obtained by the decision maker.
6 Comparative Statics
We first define comparative risk preferences in a standard way. This will help us to contrast comparative risk aversion with other comparative results.
@@ 6 Comparative Statics
$\succsim^{2}$ is at least as risk averse than $\succsim^{1}$ if whenever $1/2x\oplus 1/2y\succsim^{1}z$, then $1/2x\oplus 1/2y\succsim^{2}z$. Two decision makers are equally risk averse if each is at least as risk averse as the other.
@@
The following result shows that risk attitude is unrelated to the attitude towards verifiable events. Thus, the concern for verifiability cannot be captured by risk aversion and vice versa. We can see risk aversion as an aversion to ex-ante uncertainty of perfectly verifiable consequences and preference for verifiability as an aversion to ex-post uncertainty. Moreover, we only require identical preferences over two consequences in order to compare two decision makers with respect to their verifiable events.
@@
The result shows why critical events are at the core of preference for verifiability. If one decision maker’s set of critical events is a subset of the other decision maker’s, then it is also necessarily the case that their set of verifiable events is a subset.
In the context of our example, we may care about how inefficient a decision maker’s choice is when following an expected verifiability utility or an expected obfuscation utility. A plausible benchmark is expected utility maximization because the agent is probabilistically sophisticated over verifiable events.
@@
It is noteworthy that from observing expected verification preferences alone we cannot uniquely determine the welfare loss because the probabilities of non-verifiable events are not uniquely determined. If some consequences are not perfectly verifiable, we need to complement the preference information with probability information of non-verifiable events. However, it turns out that even if we make the assumption that a policy maker picks the “correct” $\mu$, policy interventions that change the set of verifiable events or that change firms’ behavior from obfuscation-seeking to verification-seeking do not necessarily reduce the welfare loss. The following examples show that (holding fixed the verifiable events) a decision maker acting according to expected verification utility may indeed face a larger welfare loss than a decision maker acting according to expected obfuscation utility and that increasing the set of verifiable events need not lead to a welfare improvement. This is noteworthy because this implies that calls for better verifiability of CO2 reduction and/or incentives against greenwashing need to be carefully evaluated for whether they will achieve the desired purpose.
@@
    There exist $\mu$ and decision problems $A$ and $A^{\dagger}$ such that $L_{\succsim,\mu}(A)>L_{\succsim^{\dagger},\mu}$ and $L_{\succsim^{\dagger},\mu}(A^{\dagger})>L_{\succsim,\mu}(A^{\dagger})$.
Even though we may intuitively find that behavior according to expected verification utility is normatively more appealing than that of expected obfuscation utility, the expected utility loss may be larger.
@@
Suppose in the example decision problem states $s$ and $u$ are very unlikely. In this case, an expected utility maximizer would choose RECs, the same action as a maximizer of expected obfuscation utility. A maximizer of expected verifiability utility would choose Trees and incur a welfare loss. This is because the decision maker fears being blamed for an ex-post suboptimal action that was ex-ante optimal. ∎
@@
In our example, suppose $\mathscr{V}^{\prime}=\{\{s\},\{s,t,u\}\}$ and $\mathscr{V}=\{\{s,t,u\}\}$. Let $A$ contain only the act Trees and RECs with a slight payoff increase of RECs in state $u$. Suppose the beliefs are such that $\mu(\{t\})=.99$, meaning that an expected utility maximizer would strictly prefer RECs. Under verifiable events $\mathscr{V}$, the decision maker would prefer RECs and there is zero welfare loss. In $\mathscr{V}^{\prime}$ the decision maker is able to verify that event $\{s\}$ obtains, and now prefers Trees over RECs because in state $s$ the act Trees yields a higher payoff. However, this leads to a strictly positive welfare loss despite providing more information. Together, this and the previous example show that policy decisions regarding the incentives for transparency and efficiency of carbon emission reduction are nontrivial. While transparency and verifiability are perhaps intrinsically desirable, they may come at the cost of firms choosing less efficient CO2 reduction strategies. This suggests that policies that are aimed at increasing ex-post verifiability need to be carfully examined whether the gain in transparency might be offset by a loss in efficiency — unless, ideally, the policy implements full ex-post transparency. ∎

### v2

@@ 5 Axioms
This section provides a set of axioms that characterize expected verification and obfuscation utilities. The axioms allow us to distinguish between verification and obfuscation seeking behavior and should be normatively compelling if consequences are not observable.
We assume the existence of a biseparable utility representation. Biseparable preferences allow us to identify the decision weights of individuals on events while making minimal assumptions about behavior. The axioms for biseparable utility are provided in [undefad] and for convenience restated in appendix I.
@@
Thus, in the biseparable model, decision makers’ preferences are only meaningfully restricted on binary acts. Under the assumption that certainty equivalents exist and preferences are biseparable, we can define preference averages of consequences and acts. These preference averages, first introduced in [undefae], allow us to use [undefw] style axioms without resorting to objective probabilities. While objective probability lotteries can easily be implemented in a laboratory setting, for applications in the context of greenwashing behavior, such lotteries are often unavailable.
@@
Suppose Trees and RECs refer to investing a fixed amount of money into the relevant carbon offset certificates. Suppose further that there are no returns to scale to either technology. If the decision maker is risk-neutral (in terms of the curvature of $u$) over the amount of CO2 emissions reduced, then $1/2$ Trees $\oplus 1/2$ RECs refers to the act in which the carbon emissions in every state of the world are the arithmetic average of Trees and RECs. This can be thought of as investing half the money into each technology. However, if the decision maker is not risk-neutral, this is not the case. Instead, in each state $1/2$ Trees $\oplus 1/2$ RECs would offer the certainty equivalent of a (perfectly verifiable) equal probability lottery of the reduction achieved by Trees and RECs on that state. ∎
Our first axiom, comonotonic independence, is a separability condition from [undefap] that is normatively appealing in our context.
@@
Thus, two acts are comonotonic if they agree on the ranking of states according to whether they achieve more desirable consequences. For instance, two acts are not comonotonic if one yields a better consequence in state $s$ than in $s^{\prime}$, while the other yields a worse consequence in $s$ than in $s^{\prime}$. Constant acts are comonotonic with all acts.
@@
$1/2a\oplus 1/2c\succsim 1/2b\oplus 1/2c.$ (4)
Comonotonic independence is normatively compelling in our setting because deviations from expected utility should only arise from a lack of verifiability. A preference reversal in (4) should only occur if mixing with act $c$ makes a good consequence verifiable for one act but not the other. This is ruled out when all three acts are comonotonic. If $a,b,c$ are comonotonic, then for any verifiable event, the state with the worst outcome is the same for all three acts. Thus, mixing with $c$ does not asymmetrically change what can be verified. Analogously, also what can be obfuscated does not change: the best possible state in each verifiable event will remain the same.
The following result, a corollary of Möbius inversion of the capacity [[, e.g.,]]chateauneuf_characterizations_1989, makes precise why comonotonic independence is the right starting point for modeling verifiability in decisions under uncertainty:
Corollary 1 (Expected Verifiability Representation).
The following statements are equivalent:
  1. 1.
[…]
  2. 2.
[…]
    $U(a)=$ $p\int_{E\in\mathscr{E}^{+}}\min_{s\in E}u(a(s))dp^{+}$
    $+(1-p)\int_{E\in\mathscr{E}^{-}}\max_{s\in E}-u(a(s))dp^{-}$ (5)
    where $\mathscr{E}^{+}$, $\mathscr{E}^{-}$ are disjoint subsets of $\mathscr{E}$, $p^{+}$ is a strictly positive probability measure over ex-post information sets $\mathscr{E}^{+}$ and $p^{-}$ is a strictly positive probability measure over $\mathscr{E}^{-}$ and $u:\mathscr{X}\rightarrow\mathbb{R}$ a convex valued function.
This representation is effectively Choquet Expected Utility [[, CEU;]]schmeidler_subjective_1989. That is, a CEU-maximizing decision maker can be interpreted as maximizing an expectation over possible ex-post information sets. On each information set, the decision maker faces either a verification or an obfuscating situation: if the event belongs to $\mathscr{E}^{+}$, the worst consequence on that event is taken as given; if it belongs to $\mathscr{E}^{-}$, the best consequence is taken as given. Thus, with probability $p$, the decision maker expects a verification situation—needing to prove that a good outcome occurred—with $p^{+}$ governing the distribution over what will be verifiable. With probability 1p1-p, they expect an obfuscating position, with $p^{-}$ governing the distribution over ex-post available verifiable events and a reversed ranking of consequences. Because of this reversal of preferences over consequences, we immediately see that comonotonic independence together with a fixed ordering of outcomes requires the decision maker to either be obfuscation-seeking or certification-seeking, but not both.
While the representation nests our two main models as special cases, it imposes no structure on which events belong to $\mathscr{E}^{+}$ and which to $\mathscr{E}^{-}$. In particular, disjoint events $E,F\in\mathscr{E}^{+}$ may have their union $E\cup F\in\mathscr{E}^{-}$: the decision maker pessimistically evaluates $E$ and $F$ individually yet optimistically evaluates any act whose outcome is only known to lie in $E\cup F$. This is difficult to reconcile with a coherent model of verifiability—what can be proven ex-post should be determined by the state that occurred, not by an independent draw from $p^{+}$ and $p^{-}$ that lives outside the state space. The axioms we impose next restore this coherence.
We therefore now impose axioms on the structure of what is verifiable ex-post. Our main identification tools will be sensitive events and reactive events, that help us identify what the decision maker expects to be ex-post verifiable: If a decision maker only cares about the good consequences that they can verify ex-post, they will often be unresponsive to improvements of acts on nonnull states if these improvements are not verifiable. Conversely, if on an event the decision maker dislikes a worsening of the consequence on any state, then the decision maker must expect that if the event occurs, they will be for sure in possession of some proof that the event (or a subevent) obtained. We thus define:
Definition 5 (Sensitive event).
Let $\gamma\succ\beta$. An event $E$ is sensitive if $\gamma_{E}\beta\succ\gamma_{E-F}\beta$ for all nonnull events $F\subset E$.
Thus, an event is sensitive if getting a worse outcome on any nonnull subevent decreases preference. For an expected utility maximizer, all nonnull events are sensitive. In certification utility, an event $E$ is sensitive if and only if the verifiable subsets of $E$ form a cover of EE—that is, every state in $E$ belongs to at least one verifiable subevent of $E$.
In obfuscation utility, when a stakeholder confronts the decision maker with evidence that event $E$ obtained, the decision maker responds by pointing to the best possible consequence within $E$. Hence, even a single good outcome anywhere in $E$ provides a defense—a strictly bad consequence can only be proven if every state in $E$ yields a bad outcome. As a result, the decision maker strictly benefits from any improvement on any nonnull subevent of $E$. We thus define:
Definition 6 (Reactive event).
Let $\gamma\succ\beta$. An event $E$ is reactive if $\beta_{E-F}\gamma\succ\beta_{E}\gamma$ for all nonnull events $F\subset E$.
A reactive event therefore requires only an improvement on any nonnull event and this increases the preference of the decision maker. Again, in expected utility all events are reactive. In our obfuscation utility, reactive events are the events $E$ for which the stakeholder has a cover of verifiable events in $\mathscr{V}$ that are all subsets of $E$.
For example, if $E-\{s\}$ is verifiable but $E$ itself is not, a good outcome in state $s$ never helps the decision maker. When the true state is $s$, any verifiable event the stakeholder can use must extend beyond $E$, so the decision maker can already point to a good outcome outside $E$. When the true state is in $E-\{s\}$, state $s$ does not appear in the evidence at all.
The decision maker can combine proofs of events. This generates a combined proof that the intersection of the two events has occurred. This is the standard assumption of verifiability in contract theory [undefx].
Axiom 2 (Proof Conjunction).
[…]
To be precise, sensitive events are events for which covers of potentially many proofs exist. However, if both $E$ and $F$ have such a cover of proofs, then we can form a cover for $E\cap F$ by intersecting every verifiable event in the cover of $E$ with every verifiable event in the cover of $F$.
@@
Suppose the firm can verify both $\{s,t\}$ (by presenting purchase certificates, showing certificate supply was sufficient) and $\{t,u\}$ (via an academic study showing that purchasing RECs in this market does not cause energy-market demand shifts). Proof Conjunction requires that $\{s,t\}\cap\{t,u\}=\{t\}$ is also a sensitive event. The firm can present both proofs simultaneously when state $t$ is the true state: the certificate proof of $\{s,t\}$ rules out state $u$, while the study proof of $\{t,u\}$ rules out state $s$. Together, they certify that the true state is precisely $\{t\}$, where RECs achieves 100 megatons compared to 70 for Trees. If state $t$ is sufficiently likely, this additional proof therefore makes RECs the preferred choice over Trees. ∎
The final key axiom makes precise that the decision maker is an expected utility maximizer under the condition that every consequence is verifiable. We impose this via ambiguity neutrality from the literature on (ex-ante) ambiguity attitudes.
Axiom 3 (Proof Sufficiency).
@@
This axiom states that if an event is covered by two sensitive events $E$ and $F$, then the bets on $A\cap E$, i.e., act $f$ and bets on $A\cap F$, i.e., act $g$, are ambiguity neutral, i.e., there is no motive for hedging; the mixed act $1/2f\oplus 1/2g$ is indifferent to both $f$ and $g$.
@@
Again suppose both $E=\{s,t\}$ and $F=\{t,u\}$ are sensitive. Consider first $A=\{s,t,u\}$, so $f=\gamma_{\{s,t\}}\beta$ and $g=\gamma^{\prime}_{\{t,u\}}\beta^{\prime}$. The good outcomes of both acts fall on verifiable events: the certificate proof establishes $\gamma$ for $f$ whenever $s$ or $t$ obtains, and the study proof establishes $\gamma^{\prime}$ for $g$ whenever $t$ or $u$ obtains. The consequences of both acts are fully verifiable in every state and are therefore evaluated like expected-utility acts and mixing these acts should provide no benefit.
Now consider $A=\{s,u\}$, so $f=\gamma_{\{s\}}\beta$ and $g=\gamma^{\prime}_{\{u\}}\beta^{\prime}$: the good outcomes now fall on non-verifiable states and indifference of the two acts requires $\beta\sim\beta^{\prime}$. Under $f$, the firm achieves $\gamma$ in state $s$, but the certificate proof of $\{s,t\}$ only certifies the worst outcome on $E$, which is $\beta$ in state $t$. The good outcome is not verifiable. The same applies to $g$ in state $u$: the study proof of $\{t,u\}$ certifies at most $\beta^{\prime}$, again because state $t$ yields $\beta^{\prime}$ under $g$. State $t$ is a bottleneck in both proofs. Mixing the two acts does not resolve this: state $t$ still yields a bad outcome $1/2\beta\oplus 1/2\beta^{\prime}$. ∎
5.1 Characterization of Expected Certification Utility
The above stated axioms contain necessary and sufficient conditions on preferences for an expected certification utility. More precisely, we obtain the following characterization result:
@@ 5.1 Characterization of Expected Certification Utility
Suppose $\succsim^{1}$ and $\succsim^{2}$ are expected certification utilities. Then $a\succsim^{1}b\Leftrightarrow a\succsim^{2}b$ for all $a,b\in\mathscr{A}$ if and only if:
@@
In our example, the verifiability of $\{s,t\}$ and $\{s,t,u\}$ can be inferred from behavior of the decision maker and need not be exogenously given. Moreover, the subjective probabilities of $\{s,t\}$ and $\{u\}$ are uniquely determined from behavior. ∎
Effectively, when preferences over acts are identical then if one decision maker can prove that a certain event obtains, the other decision maker is able to prove that this event obtains or even able to prove that a strict subset of this event obtains. Thus, it could be the case that decision maker 1 is in possession of proofs that $E,F,E\cup F,E\cap F$ obtain (whenever they actually obtain) while decision maker 2 is only able to prove that $E,F,E\cap F$ obtain. In terms of behavior, we would not see a difference because the decision maker will always want to provide the most precise proof possible. Thus, we cannot identify each proof uniquely but only the collection of all proofs by closure under unions, the sensitive events. Of course, this corresponds to a unique minimal cover of the state space with verifiable events.
@@ 5.2 Characterization of Expected Obfuscation Utility
We also obtain a dual theorem that characterizes obfuscation utility. Since an expected obfuscation utility is equivalent to an expected certification utility with a reverse order of preferences, supermodularity is replaced by submodularity and proof sufficiency holds with respect to reactive instead of sensitive events.
@@
Naturally, the expected obfuscation utility has the same uniqueness properties as expected certification utility. Expected obfuscation utility can be interpreted as the decision maker minimizing the utility of a stakeholder with expected certification utility who has a reverse order of preference of consequences. Similarly, expected certification utility can be seen as minimizing the utility of a stakeholder who wants to nitpick, i.e., who wants to point at the possibility that a bad consequence may have been obtained by the decision maker.
6 Model Properties
In this section we formally derive the connection of verification and obfuscation preferences to ambiguity attitudes and comparative statics of the model.
Definition 7 (Ambiguity Attitude).
A decision maker is ambiguity averse if for any two acts $f,g$ with $f\sim g$, we have $\frac{1}{2}f+\frac{1}{2}g\succsim f$. A decision maker is ambiguity seeking if for any two acts $f,g$ with $f\sim g$, we have $\frac{1}{2}f+\frac{1}{2}g\precsim f$.
We did not impose any preference for or against ambiguity in our axioms but merely that for certain binary acts the decision maker is ambiguity neutral. Nonetheless, it turns out that the axioms are sufficiently strong to guarantee ambiguity averse behavior for verification preferences and ambiguity seeking behavior for obfuscation preferences:
Corollary 2 (Ambiguity Attitude).
A decision maker with an expected verification (obfuscation) utility is ambiguity averse (seeking).
Thus, our expected certification utility model is a special case of both Choquet expected utility with a supermodular capacity [undefap] and maxmin expected utility [undefag].
For performing comparative statics, we first define comparative risk preferences in the standard way. This will help us to contrast comparative risk aversion with other comparative results.
@@ 6 Model Properties
$\succsim^{2}$ is at least as risk averse as $\succsim^{1}$ if whenever $1/2x\oplus 1/2y\succsim^{1}z$, then $1/2x\oplus 1/2y\succsim^{2}z$. Two decision makers are equally risk averse if each is at least as risk averse as the other.
@@
The following result shows that risk attitude is independent of the attitude towards verifiable events. Thus, the concern for verifiability cannot be captured by risk aversion and vice versa. We can see risk aversion as an aversion to ex-ante uncertainty of perfectly verifiable consequences and preference for verifiability as an aversion to ex-post uncertainty. Moreover, we only require identical preferences over two consequences in order to compare two decision makers with respect to their verifiable events.
@@
The result shows why sensitive events are at the core of preference for verifiability. If one decision maker’s set of sensitive events is a subset of the other decision maker’s, then it is also necessarily the case that their set of verifiable events is a subset.
In the context of our example, we may care about how inefficient a decision maker’s choice is when following an expected certification utility or an expected obfuscation utility. A plausible benchmark is expected utility maximization because the agent is probabilistically sophisticated over verifiable events.
@@
It is noteworthy that from observing expected verification preferences alone we cannot uniquely determine the welfare loss because the probabilities of non-verifiable events are not uniquely determined from behavior. If some consequences are not perfectly verifiable, we need to complement the preference information with probability information of non-verifiable events.
Given a fixed probability distribution over states it may be expected that increases in transparency generally decrease the welfare loss and align behavior closer with expected utility maximization or that it is more desirable for a decision maker to be certification-seeking than obfuscation-seeking. However, policy interventions that change the set of verifiable events or that change firms’ behavior from obfuscation-seeking to certification-seeking turn out to not necessarily reduce the welfare loss given a fixed $\mu$. The following examples show that (holding fixed the verifiable events) a decision maker acting according to expected certification utility may indeed face a larger welfare loss than a decision maker acting according to expected obfuscation utility and that increasing the set of verifiable events need not lead to a welfare improvement. This is noteworthy because this implies that calls for better verifiability of CO2 reduction and/or incentives against greenwashing need to be carefully evaluated for whether they will achieve the desired purpose.
@@
    There exist $\mu$ and decision problems $A$ and $A^{\dagger}$ such that $L_{\succsim,\mu}(A)>L_{\succsim^{\dagger},\mu}(A)$ and $L_{\succsim^{\dagger},\mu}(A^{\dagger})>L_{\succsim,\mu}(A^{\dagger})$.
Even though we may intuitively find that behavior according to expected certification utility is normatively more appealing than that of expected obfuscation utility, the expected utility loss may be larger or smaller, depending on the available actions.
@@
Suppose in the example decision problem states $s$ and $u$ are very unlikely. In this case, an expected utility maximizer would choose RECs, the same action as a maximizer of expected obfuscation utility. A maximizer of expected certification utility would choose Trees and incur a welfare loss. This is because the decision maker fears being blamed for an ex-post suboptimal action that was ex-ante optimal. ∎
@@
In our example, suppose $\mathscr{V}^{\prime}=\{\{s\},\{s,t,u\}\}$ and $\mathscr{V}=\{\{s,t,u\}\}$. Let $A$ contain only the act Trees and RECs with a slight payoff increase of RECs in state $u$. Suppose the beliefs are such that $\mu(\{t\})=.99$, meaning that an expected utility maximizer would strictly prefer RECs. Under verifiable events $\mathscr{V}$, the decision maker would prefer RECs and there is zero welfare loss. In $\mathscr{V}^{\prime}$ the decision maker is able to verify that event $\{s\}$ obtains, and now prefers Trees over RECs because in state $s$ the act Trees yields a higher payoff. However, this leads to a strictly greater welfare loss despite providing more information. Together, this and the previous example show that policy decisions regarding the incentives for transparency and efficiency of carbon emission reduction are nontrivial. While transparency and verifiability are perhaps intrinsically desirable, they may come at the cost of firms choosing less efficient CO2 reduction strategies, see [undefan] for a study on afforestation. This suggests that policies that are aimed at increasing ex-post verifiability need to be carefully examined whether the gain in transparency might be offset by a loss in efficiency — unless, ideally, the policy implements full ex-post transparency. ∎
From these ex-ante counterintuitive results it follows that policies aimed at increasing but not fully implementing transparency or inducing certification seeking behavior need to take into account the available actions of the decision makers. This result already holds without addressing limiting factors of this welfare analysis; in general, the decision maker’s preferences may not perfectly align with society’s and costs of different policies are not explicitly modeled.